The tensile strength Y of a certain material is such that Y=e^{^{x}} where X is normally distributed with a mean of \mu =10 and a variance of \sigma ^{2} =1 (Y is said to follow a log-normal distribution. Compute each of the following:

a. P(Y<7,000) [hint: (lnY-\mu)/\sigma~N(0,1)]

b. P(10,000<Y<20,000)

c. P(Y>15,000)
Answer #1

a. P(Y<7,000)

In(y) _ μ P( In(7000) _ 10

= P(Z< -1.1463) = 0.1258

b. P(10,000<Y<20,000)

= P(\frac{ln(10000)-10}{1} < \frac{ln(Y) - \mu}{\sigma} < \frac{ln(20000)-10}{1})

= P(-0.7897<Z< -0.0965) = 0.2467

c. P(Y>15,000)

= P(\frac{ln(Y) - \mu}{\sigma} > \frac{ln(15000)-10}{1})

= P(Z> -0.3842) = 0.6496